# IEEE Copyright Notice

### v2

@@ Abstract
Data-driven learning of soft continuum robot (SCR) dynamics from high-dimensional observations offers flexibility but often lacks physical interpretability, while model-based approaches require prior knowledge and can be computationally expensive. We bridge this gap by introducing (1) the Attention Broadcast Decoder (ABCD), a plug-and-play module for autoencoder-based latent dynamics learning that generates pixel-accurate attention maps localizing each latent dimension’s contribution while filtering static backgrounds. (2) By coupling these attention maps to 2D oscillator networks, we enable direct on-image visualization of learned dynamics (masses, stiffness, and forces) without prior knowledge. We validate our approach on single- and double-segment SCRs, demonstrating that ABCD-based models significantly improve multi-step prediction accuracy: 5.7× error reduction for Koopman operators and 3.5× for oscillator networks on the two-segment robot. The learned oscillator network autonomously discovers a chain structure of oscillators. Unlike standard methods, ABCD models enable smooth latent space extrapolation beyond training data. This fully data-driven approach yields compact, physically interpretable models suitable for control applications.
@@ I Introduction
Figure 1: Main contributions of this study: (1) A newly proposed attention broadcast decoder (ABCD) is used plug-and-play for autoencoder-based Koopman and oscillator dynamics learning of a soft continuum robot. (2) Attention maps inside the ABCD are coupled to 2D oscillator dynamics so that the oscillator network can be visualized (here, masses, stiffnesses and actuation forces) on the original, reconstructed, or predicted future images.
The ability to learn dynamical systems directly from high-dimensional observations has recently transformed how we model and control complex robots. For example, Dynamic Mode Decomposition (DMD), Koopman operator theory, and deep learning-based embeddings enable data-driven discovery of latent dynamics without explicit modeling [brunton2022modern, lusch2018deep, li2020learning, takeishi_learning_2017]. Koopman operator methods offer the advantage of providing a linear representation of nonlinear dynamics, which facilitates prediction and control using linear methods [brunton2022modern].
Using (coupled) oscillator networks to describe latent dynamics, on the other hand, has been shown to be a versatile, expressive low-order time sequence modeling approach that offers potential (though not often used) for interpretability [lanthaler2023neural]. Learning dynamics directly from video data is attractive for both methods, as it requires minimal setup (no markers, trackers, or complex post-processing), no prior knowledge, and allows end-to-end learning. Building on this idea, Stölzle and Della Santina demonstrated that input-to-state stable coupled oscillator networks (CONs) can learn latent-space dynamics of soft continuum robots (SCRs) from generated video data and support model-based control [stoelzle2024input].
SCRs are an especially challenging research platform due to their infinite-dimensional, highly nonlinear mechanics [laschi_learningbased_2023]. Even if accurate models exist, they are often too computationally expensive for real-time control. As a result, data-driven dynamics learning has become an increasingly important direction in soft robotics. Recent reviews highlight the growing trend toward learning-based dynamics models for robotic manipulation [ai2025review].
Overall, a trade-off exists between interpretability, performance, and modeling accuracy. Data-driven methods provide flexibility and simplicity but often lack interpretability and physical meaning. Model-driven approaches are interpretable and physically meaningful but require extensive prior modeling and are computationally heavy. Hybrid methods offer a promising compromise but still rely on manually designed physics components. To the best of our knowledge, no existing approach is fully data-driven yet yields a compact, physically interpretable low-parameter model suitable for control of soft continuum robots.
@@
    proposing a novel Attention BroadCast Decoder (ABCD) for autoencoder-based latent dynamics learning from image sequences (video). It assigns pixel-accurate attention maps for each latent, indicating the location of the latents in the image while fully filtering out the static background.
@@
    showing how the attention maps of the ABCD can be coupled with oscillator network learning to enable physically meaningful on-image visualization of a 2D oscillator network. This achieves unprecedented levels of interpretability of learned dynamics and identification of low-order dynamics without prior knowledge.
This denotes a novel approach for deriving an SCR model from video data without prior knowledge, exhibiting a level of physical structure and interpretability similar to manual models.
@@ II-A Hybrid and Data-Driven Modeling of SCRs
Koopman-based models lift the nonlinear dynamics of soft continuum robots to a linear control space enabling real-time model predictive control (MPC) [bruder_datadriven_2021, shi2023koopman]. Haggerty et al. [haggerty2023control] showed that Koopman methods can even be used to control highly dynamic motions of soft continuum robots. Ristich et al. [ristich_physicsinformed_2025b] introduced a hybrid Koopman approach that improves simulation accuracy from small datasets by additionally using phase-space information from a Cosserat rod model.
Physics-informed neural networks (PINNs) combine physical priors and data for highly flexible yet structured learning. Liu et al. [liu_physicsinformed_2024] use Lagrangian and Hamiltonian neural networks to control a tendon-driven SCR in simulation; however, real-robot results remain limited to quasi-static predictions due to the simplified underlying physics model. Licher et al. [licher2025adaptive] train a domain-decoupled PINN [krauss_domaindecoupled_2024b] on a full Cosserat rod model and utilize it for adaptive MPC and state estimation.
Reduced-order modeling (ROM) methods also play a significant role in SCR modeling. Alkayas et al. [alkayas2025structure] propose a structure-preserving autoencoder for fast simulation and control; in related work [alkayas2025soft], an optimal strain parameterization based on proper orthogonal decomposition is introduced. These methods reduce system dimensionality while retaining essential physical structure.
Each of these methods has limitations: Koopman-based models do not offer a direct mechanistic interpretation, PINNs require manual derivation of physics models, and ROMs depend on accurate high-fidelity simulations that must also be derived by hand.
@@ II-B Vision-Based and Image-Driven Modeling of Soft Robots
Recent progress in vision-based learning highlights the potential of image-only sensing for soft robots. Zheng et al. [zheng2024vision] achieved online estimation of key points for multiple SCRs using stereo images and convolutional neural networks (CNNs). Rong and Gu [rong2024vision] used CNNs for real-time shape estimation of self-occluding soft parallel robots. Almanzor et al. [almanzor_static_2023] utilized a CNN as a deep visual inverse-kinematics model for static shape control.
To also consider dynamics, Monteiro et al. [monteiro2024visuo] developed a visuo-dynamic self-modeling approach using recurrent networks, which learns from actuation inputs and video data to predict and control the motion of soft robots. A more interpretable hybrid model is presented by Valadas et al. [valadas_learning_2025a], who learn a piecewise-constant strain model from video data and integrate it into a model-based controller.
None of these vision-based methods directly yields a compact and physically interpretable dynamical model. The approach of Stölzle and Della Santina [stoelzle2024input] comes closest, yet their oscillators are not linked to the robot’s physical configuration or structural elements, and the method is only demonstrated on simulated SCRs.
@@ III-A Autoencoder-Based Latent Dynamics Learning
We implement an autoencoder-based latent dynamics learning approach for either Koopman or oscillator network learning based on and modified from [lusch2018deep, stoelzle2024input].
@@
with the encoder network $\varphi$. We then predict the next latent coordinate
$\bm{z}^{(i+1)}=f_{\text{dyn}}(\bm{z}^{(i)},\bm{\dot{z}}^{(i)},\bm{u}^{(i)})$ (2)
after time interval $\Delta t$ with the dynamical model $f_{\text{dyn}}$, with $\varphi$ and $\varphi^{-1}$ being approximated by neural networks. We can then retrieve the predicted next image
$\hat{\bm{o}}^{(i+1)}=\varphi^{-1}(\bm{z}^{(i+1)})$ (3)
with the decoder network $\varphi^{-1}$. This approach is similar to [lusch2018deep].
The latent velocities $\bm{\dot{z}}^{(i)}$ together with the latent coordinates $\bm{z}^{(i)}$ form the latent state. The latent velocities are obtained by mapping observation-space velocities to latent space via the encoder’s Jacobian after [stoelzle2024input]. Observation velocities are approximated using central finite differences which are then mapped to latent velocities using forward-mode automatic differentiation (AD)
@@
We use a $\beta$-VAE [higgins2017beta] for the encoder and decoder networks. Following [lusch2018deep], the model is trained end-to-end with the basic loss
@@
Symplectic Euler integration is used during training for stability. For oscillator networks, we augment $\mathcal{L}_{\text{basic}}$ with a steady-state loss $\lambda_{\text{s}}\mathcal{L}_{\text{s}}$ to enforce $\bm{z}^{(i)}=\bm{z}_{0}$ and $\bm{\dot{z}}^{(i)}=\bm{0}$ at equilibrium, encouraging physically consistent steady-state dynamics.
@@ III-B1 Architecture
We introduce the Attention Broadcast Decoder (ABCD) as a plug-and-play module for autoencoder-based latent dynamics learning. It is inspired by the spatial broadcast decoder of Watters et al. [watters2019spatial], originally proposed for disentangled representation learning. As shown in Fig. 2(a), the ABCD takes latent coordinates $\bm{z}^{(i)}$ and produces reconstructed images and pixel-accurate attention maps. It shall be noted that while the ABCD is used as a replacement for a standard CNN decoder, the ABCD implicitly affects the structure of the latent state and the dynamics model.
@@
are calculated by concatenating the logits with a constant background logit $b\equiv\ell^{(i)}_{k+1}(x,y)$ (a hyperparameter) and normalizing via softmax, resulting in attention maps $\bm{a}_{j}^{(i)}\in\mathbb{R}^{h\times w}$ for $j=1,\ldots,k+1$, including the background attention map $\bm{a}_{b}^{(i)}$. Optionally, adaptive Gumbel noise—controlled by a gumbel noise strength parameter—can be added to the logits before the softmax to encourage sharper, more competitive attention during training. Separately, each scalar latent $z_{j}^{(i)}$ is expanded to extended latents
@@
via a linear layer of parameters $\bm{W}_{j}$ and $\bm{b}_{j}$, mapping to feature width $n_{f}$ and then broadcasted to image resolution, yielding $\overline{\bm{z}}_{j}^{(i)}\in\mathbb{R}^{n_{f}\times h\times w}$. Additionally, learnable spatial background features $\overline{\bm{z}}_{b}\in\mathbb{R}^{n_{f}\times h\times w}$ provide an auxiliary expanded feature vector for capturing static background content. The attention-weighted latent coordinate maps are then computed by
@@
where $\odot$ denotes element-wise multiplication. These are concatenated with coordinates and processed by four $1{\times}1$ CNN layers to construct the output image. Unlike the original spatial broadcast decoder [watters2019spatial] with the latent state uniformly broadcasted on the image resolution before final encoding, our approach makes individual latents compete for reconstruction of every output image pixel via the attention maps, i.e., yielding attention maps for each latent and background that show its relative contribution to the reconstruction of each pixel. Additionally, the introduced learnable background features allow to fully decouple the static background reconstruction from the latent state.
@@ III-B2 Attention consistency loss
to regularize the attention dynamics, where $\dot{a}_{j}^{(i)}(x,y)$ is the time derivative (velocity) of the attention map for latent $j$ at spatial location $(x,y)$, computed via a Jacobian-vector product that propagates the observation velocity through the attention mechanism. Further, $\dot{\bar{\bm{o}}}^{(i)}(x,y)$ is the normalized velocity of the observation at pixel $(x,y)$. This loss penalizes changes in the attention map at locations where the underlying image does not change, thereby encouraging the model to assign dynamic attention only to non-static regions. This creates a sharper transition between background and attention maps and stronger aligns attention dynamics with image dynamics.
III-B3 2D oscillator networks and attention coupling
For oscillator network learning when used with the ABCD, we introduce an additional modification that interprets the learned dynamics as a 2D oscillator network coupled to the attention maps to enable on-image visualization. We group consecutive pairs of latent dimensions into $n=k/2$ two-dimensional oscillators, where oscillator $l\in\{1,\ldots,n\}$ has latent-space position $\bm{q}_{l}^{(i)}=[z_{2l-1}^{(i)},z_{2l}^{(i)}]^{\top}\in\mathbb{R}^{2}$ and velocity $\dot{\bm{q}}_{l}^{(i)}=[\dot{z}_{2l-1}^{(i)},\dot{z}_{2l}^{(i)}]^{\top}\in\mathbb{R}^{2}$. We modify the ABCD’s attention processor to generate one attention map $\bm{a}_{l}^{(i)}\in\mathbb{R}^{h\times w}$ per oscillator (instead of per individual latent dimension). Correspondingly, the mass matrix $\bm{M}$ is constrained such that $M_{2l-1,2l-1}=M_{2l,2l}$ for all $l$, ensuring equal mass for both dimensions of each 2D oscillator, while the damping and stiffness matrices $\bm{D}$ and $\bm{K}$ remain unconstrained. To identify oscillator positions into image space, we compute the center-of-mass (COM) of the squared attention maps:
@@ III-C Soft Continuum Robot Datasets
For this study the dynamic motion of one SCR with a single segment and one SCR consisting of two sequential connected by a rigid connector are investigated. The segments are soft pneumatic actuators with an outer diameter of $42.4\text{\,}\mathrm{m}\mathrm{m}$ and a length of $130\text{\,}\mathrm{m}\mathrm{m}$. They entail three fiber-reinforced pneumatic chambers arranged at $120\text{\,}\mathrm{\SIUnitSymbolDegree}$ intervals around an inner channel. The actuator’s main body is molded from EcoFlex 00-50 silicone (Smooth-On), while the base and tip caps are fabricated from Dragon Skin 30 silicone (Smooth-On). These actuators, in general, can generate spatial motion. However, only planar motion can be observed by a single camera. To ensure that the actuator only moves in a plane perpendicular to the camera axis, two of the actuator’s chambers are connected to the same valve and therefore always hold the same pressure. We used 5/3 proportional valves (Enfield, LS-V05s) and absolute pressure sensors (First Sensor, 142BC30A-PCB), to control at 1 kHz using a PID controller implemented in Simulink. Further information about the actuator design and manufacturing as well as the pressure control setup are given in [Anonymous_2021]. The video is recorded using a Sony Alpha 7 IV camera at $120\text{\,}\mathrm{f}\mathrm{p}\mathrm{s}$ in Full HD.

### v3

@@ Abstract
Learning soft continuum robot (SCR) dynamics from video offers flexibility but existing methods lack interpretability or rely on prior assumptions. Model-based approaches require prior knowledge and manual design. We bridge this gap by introducing: (1) The Attention Broadcast Decoder (ABCD), a plug-and-play module for autoencoder-based latent dynamics learning that generates pixel-accurate attention maps localizing each latent dimension’s contribution while filtering static backgrounds, enabling visual interpretability via spatially grounded latents and on-image overlays. (2) Visual Oscillator Networks (VONs), a 2D latent oscillator network coupled to ABCD attention maps for on-image visualization of learned masses, coupling stiffness, and forces, enabling mechanical interpretability. We validate our approach on single- and double-segment SCRs, demonstrating that ABCD-based models significantly improve multi-step prediction accuracy with 5.8× error reduction for Koopman operators and 3.5× for oscillator networks on a two-segment robot. VONs autonomously discover a chain structure of oscillators. This fully data-driven approach yields compact, mechanically interpretable models with potential relevance for future control applications.
@@ I Introduction
Figure 1: Main contributions of this study: (1) A newly proposed attention broadcast decoder (ABCD) is used plug-and-play for autoencoder-based Koopman and oscillator dynamics learning of a soft continuum robot. (2) Attention maps inside the ABCD are coupled to a 2D latent oscillators network to achieve VONs, so that masses, coupling stiffness, stiffness forces, inertial forces, and actuation forces can be visualized on the original, reconstructed, or predicted future images.
The ability to learn dynamical systems directly from high-dimensional observations has recently transformed how we model and control complex robots. Koopman operator theory and deep learning-based embeddings enable data-driven discovery of latent dynamics without explicit modeling [6, 15, 12, 22]. Koopman operator methods provide a linear representation of nonlinear dynamics, facilitating prediction and control using linear methods [6].
Using (coupled) oscillator networks to describe latent dynamics is a versatile low-order time sequence modeling approach with interpretability potential [10]. Learning dynamics directly from video data requires minimal setup and enables end-to-end learning. Building on this idea, Stölzle and Della Santina showed that input-to-state stable, coupled oscillator networks (CONs) can learn latent-space dynamics of SCRs from simulated video and support model-based control [21].
SCRs are an especially challenging research platform due to their infinite-dimensional, highly nonlinear mechanics [11]. Even if accurate models exist, they are often too computationally expensive for real-time control. As a result, data-driven dynamics learning has become an increasingly important direction in soft robotics. Recent reviews highlight the growing trend toward learning-based dynamics models for robotic manipulation [1].
Overall, a trade-off exists between interpretability, performance, and modeling accuracy. Data-driven methods provide flexibility and simplicity but often lack interpretability and physical meaning. Model-driven approaches are interpretable and physically meaningful but require extensive prior modeling and are computationally heavy. Hybrid methods offer a promising compromise but still rely on manually designed physics components. To the best of our knowledge, no existing approach is fully data-driven yet yields a compact, mechanically interpretable low-parameter model suitable for control of SCRs.
@@
    proposing a novel Attention BroadCast Decoder (ABCD) for autoencoder-based latent dynamics learning from image sequences (video). This enables visual interpretability by assigning pixel-accurate attention maps for each latent, indicating its location in the image while filtering out the static background.
@@
    demonstrating how the attention maps produced by the ABCD can be coupled with a latent oscillator network to form Visual Oscillator Networks (VONs), enabling on-image depiction of explicit oscillator models whose masses, coupling stiffnesses, and forces are mechanically interpretable, in that their configuration and resulting motion corresponds to the real system. This achieves unprecedented levels of interpretability of learned dynamics and identification of low-order dynamics without prior knowledge.
This denotes a novel approach for deriving an SCR model from video data without prior knowledge, exhibiting a level of mechanical structure and interpretability similar to manual models.
@@ II-A Hybrid and Data-Driven Modeling of SCRs
Koopman-based models can lift the nonlinear dynamics of SCRs to a linear control space enabling real-time model predictive control (MPC) [5, 20]. Haggerty et al. [7] showed that Koopman methods can even be used to control highly dynamic motions of SCRs. Ristich et al. [18] introduced a hybrid Koopman approach that improves simulation accuracy from small datasets by additionally using phase-space information from a Cosserat rod model.
Physics-informed neural networks (PINNs) combine physical priors and data for highly flexible yet structured learning. Liu et al. [14] use Lagrangian and Hamiltonian neural networks to control a tendon-driven SCR in simulation. However, real-robot results remain limited to quasi-static predictions due to the simplified underlying physics model. Licher et al. [13] train a domain-decoupled PINN [9] on a full Cosserat rod model and utilize it for adaptive MPC and state estimation.
Reduced-order modeling (ROM) methods also play a significant role in SCR modeling. Alkayas et al. [3] propose a structure-preserving autoencoder for fast simulation and control, with related work on optimal strain parameterization [2].
Each of these methods has limitations: Koopman-based models do not offer a direct mechanical interpretation, PINNs require manual derivation of physics models, and ROMs depend on accurate high-fidelity simulations that must also be derived by hand.
@@ II-B Vision-Based and Image-Driven Modeling of Soft Robots
Recent progress in vision-based learning highlights the potential of image-only sensing for soft robots. Zheng et al. [25] estimate key points for SCRs from stereo images, Rong and Gu [19] perform real-time shape estimation under self-occlusion, and Almanzor et al. [4] learn deep visual inverse kinematics for static shape control.
To also consider dynamics, Monteiro et al. [16] developed a visuo-dynamic self-modeling approach using recurrent networks, which learns from actuation inputs and video data to predict and control the motion of soft robots. A more interpretable hybrid model is presented by Valadas et al. [23], who learn a piecewise-constant strain model from video data and integrate it into a model-based controller.
None of these vision-based methods yields a compact, mechanically interpretable dynamical model without prior system assumptions (e.g., an explicit strain/curvature model and shape extraction). In contrast, our method learns dynamics end-to-end from raw images and provides visual and mechanical interpretability via attention-localized latents and VON visualizations. The approach of Stölzle and Della Santina [21] comes closest, yet their oscillators are not linked to the robot’s physical configuration or structural elements, and the method is only demonstrated on simulated SCRs.
@@ III-A Autoencoder-Based Latent Dynamics Learning
We implement an autoencoder-based latent dynamics learning approach for either Koopman or oscillator networks based on and modified from [15, 21].
@@
with the encoder network $\varphi$. The latent state is defined as $\bm{\xi}^{(i)}=[\bm{z}^{(i)\top},\bm{\dot{z}}^{(i)\top}]^{\top}\in\mathbb{R}^{2k}$. We then predict the next latent state
$\hat{\bm{\xi}}^{(i+1)}=f_{\text{dyn}}(\bm{\xi}^{(i)},\bm{u}^{(i)})$ (2)
after time interval $\Delta t$ with the dynamical model $f_{\text{dyn}}$. The predicted latent coordinate $\hat{\bm{z}}^{(i+1)}$ is given by the first $k$ components of $\hat{\bm{\xi}}^{(i+1)}$. With $\varphi$ and $\varphi^{-1}$ being approximated by neural networks, we can then retrieve the predicted next image
$\hat{\bm{o}}^{(i+1)}=\varphi^{-1}(\hat{\bm{z}}^{(i+1)})$ (3)
with the decoder network $\varphi^{-1}$. This approach is similar to [15].
The latent velocities $\bm{\dot{z}}^{(i)}$ are obtained by mapping observation-space velocities to latent space via the encoder’s Jacobian after [21]. Observation velocities are approximated using central finite differences which are then mapped to latent velocities using forward-mode automatic differentiation (AD)
@@
We use a $\beta$-VAE [8] for the encoder and decoder networks. Following [15], the model is trained end-to-end with the basic loss
@@
Here, $\bm{z}_{0}$ is a learned equilibrium (rest) latent coordinate and can be non-zero. Symplectic Euler integration is used during training for stability. In our implementation, the control input $\bm{u}^{(i)}$ is augmented with four delayed actuation inputs. For oscillator networks, we augment $\mathcal{L}_{\text{basic}}$ with a rest-state loss $\lambda_{\text{r}}\mathcal{L}_{\text{r}}$ to enforce $\bm{z}^{(i)}=\bm{z}_{0}$ and $\bm{\dot{z}}^{(i)}=\bm{0}$ at equilibrium, encouraging physically consistent rest-state dynamics.
@@ III-B1 Architecture
We introduce the Attention Broadcast Decoder (ABCD) as a plug-and-play module for autoencoder-based latent dynamics learning. It is inspired by the spatial broadcast decoder of Watters et al. [24]. As shown in Fig. 2(a), the ABCD takes latent coordinates $\bm{z}^{(i)}$ and produces reconstructed images and pixel-accurate attention maps. While the ABCD replaces a standard convolutional neural network (CNN) decoder, it implicitly affects the latent-state structure and the learned dynamics.
@@
are calculated by concatenating the logits with a constant background logit $b\equiv\ell^{(i)}_{k+1}(x,y)$ (a hyperparameter) and normalizing via softmax, resulting in attention maps $\bm{a}_{j}^{(i)}\in\mathbb{R}^{h\times w}$ for $j=1,\ldots,k+1$, including the background attention map $\bm{a}_{b}^{(i)}$. Separately, each scalar latent $z_{j}^{(i)}$ is expanded to extended latents
@@
via a linear layer of parameters $\bm{W}_{j}$ and $\bm{b}_{j}$ and broadcasted to image resolution. Additionally, learnable spatial background features $\overline{\bm{z}}_{b}$ provide an auxiliary feature map for capturing static background content. The attention-weighted latent coordinate maps are then computed by
@@
where $\odot$ denotes element-wise multiplication. These are concatenated with coordinates and processed by four $1{\times}1$ CNN layers to construct the output image. Unlike spatial broadcast decoders [24], the ABCD makes individual latents compete for reconstructing each pixel via the attention maps and decouples static background reconstruction from the latent state via learnable background features.
@@ III-B2 Attention consistency loss
to regularize the attention dynamics, where $\dot{a}_{j}^{(i)}(x,y)$ is the time derivative of the attention map for latent $j$ at spatial location $(x,y)$ and $\dot{\bar{\bm{o}}}^{(i)}(x,y)$ is the normalized observation velocity at pixel $(x,y)$. This loss penalizes attention changes at pixels without image motion, encouraging sharper separation between background and dynamic regions.
III-B3 Visual Oscillator Networks (VONs)
For VONs, we introduce an additional modification that interprets the learned dynamics as a VON coupled to the attention maps to enable on-image visualization. We group consecutive pairs of latent dimensions into $n=k/2$ two-dimensional oscillators, where oscillator $l\in\{1,\ldots,n\}$ has latent-space position $\bm{q}_{l}^{(i)}=[z_{2l-1}^{(i)},z_{2l}^{(i)}]^{\top}\in\mathbb{R}^{2}$ and velocity $\dot{\bm{q}}_{l}^{(i)}=[\dot{z}_{2l-1}^{(i)},\dot{z}_{2l}^{(i)}]^{\top}\in\mathbb{R}^{2}$. We modify the ABCD’s attention processor to generate one attention map $\bm{a}_{l}^{(i)}\in\mathbb{R}^{h\times w}$ per oscillator (instead of per individual latent dimension). Correspondingly, the mass matrix $\bm{M}$ is constrained such that $M_{2l-1,2l-1}=M_{2l,2l}$ for all $l$, ensuring equal mass for both dimensions of each 2D oscillator, while the damping and stiffness matrices $\bm{D}$ and $\bm{K}$ remain unconstrained. To identify oscillator positions into image space, we compute the center-of-mass (COM) of the squared attention maps:
@@ III-B3 Visual Oscillator Networks (VONs)
To visualize forces in image space after training, we map latent forces to the image plane using the local, state-dependent Jacobian $J_{l}^{(i)}=\frac{\partial p_{l}^{(i)}}{\partial q_{l}^{(i)}}$, which is analytically derived via automatic differentiation through the ABCD’s latent-to-image mapping.
@@ III-C Soft Continuum Robot Datasets
We investigate one single-segment SCR and one two-segment SCR with a rigid connector ($21\text{\,}\mathrm{m}\mathrm{m}$ long). The segments are soft pneumatic actuators casted from silicone with an outer diameter of $42.4\text{\,}\mathrm{m}\mathrm{m}$ and a length of $130\text{\,}\mathrm{m}\mathrm{m}$ and entail three fiber-reinforced chambers arranged at $120\text{\,}\mathrm{\SIUnitSymbolDegree}$ intervals around an inner channel. While the actuators can in principle move in 3D, we enforce planar motion observable by a single camera by pressure-coupling two chambers per actuator. Pressures are controlled at 1 kHz using a PID controller. Video is recorded at $120\text{\,}\mathrm{f}\mathrm{p}\mathrm{s}$ in Full HD. Further hardware and manufacturing details are provided in [17].